If you buy a call option on a Birr 200,000 bond futures contract with an exercise price of 220 and the
price of the Treasury bond is 222 at expiration, is the contract in the money, out of the money, or at
the money? What is your profit or loss on the contract if the premium was Birr 3000?
1
Expert's answer
2020-08-17T16:12:50-0400

But in my knowledge it is in the money.

222-220=2

200,000*2=400,000

2 premium=6000

or he borrowed 200,000-3000=197000

Profit=400,000-197,000=203,000

Profit=203,000